A mango is dropped from a height h.
a. If dropped from a height of 2 h, would the
impact speed double?
b. Would the air time double when dropped from
a height of 2h ?
1
Expert's answer
2020-09-21T08:13:20-0400

(a) The law of conservation of energy says

"mgh=\frac{mv^2}{2}"

The impact speed

"v=\sqrt{2gh}\sim\sqrt{h}"

Hence, if we dropped a mango from a height of "2h" the impact speed increases by factor "\sqrt{2}."

Answer: No

(b) The air time of motion

"t=\sqrt{\frac{2h}{g}}\sim\sqrt{h}"

Hence, if we dropped a mango from a height of "2h" the time of motion increases by factor "\sqrt{2}."

Answer: No